OpenStudy (sleepyjess):

One question I know I completely bombed from my final today, anyone help me figure it out please?

OpenStudy (sleepyjess):

It was Solve for x: \(3^{x-2} = 6\)

OpenStudy (sleepyjess):

I know I need to take the log of both sides, giving (x-2)log 3 = log 6, but was stuck after that

OpenStudy (bmk614):

No. You actually just put the equation in logarithmic form. Which would be \[\log_{3}6=x-2 \]

OpenStudy (bmk614):

You would then solve for x.

OpenStudy (sleepyjess):

How would I solve for x from that?

OpenStudy (bmk614):

\[\log_{3}6+2=x \]

OpenStudy (bmk614):

That should be your answer.

OpenStudy (bmk614):

To get it simplified anymore, you would need to plug it into a calculator.

OpenStudy (peachpi):

you weren't wrong with what you did. (x-2)log 3 = log 6 divide by log 3 x - 2 = (log 6)/(log 3) add 2 x = 2 + (log 6)/(log 3) which is the same as what @bmk614 got once you apply the change of base formula

OpenStudy (sleepyjess):

I was trying to distribute the x-2 and work from x log 3 - 2 log 3 = log 6, I see what I did wrong

OpenStudy (peachpi):

yeah, that's not wrong either. x log 3 - 2 log 3 = log 6 x log 3 = log 6 + 2 log 3 x = (log 6 + 2 log 3)/log 3 x = (log 6)/(log 3) + 2